What is the BMI for an individual who weighs 162 lbs and stand 65 inches tall? Is this person obese or Overweight or Underweight?
Answer #1

Body mass index (BMI) is a measure of body fat based on height and weight. It is calculated as your weight (kilograms) divided by the square of your height (in meters)

[BMI = Kg/M2]

According to the question, the weight of the individual is 162lbs = 73.48196394 kg

the height the individual is 65inches = 1.651 m

BMI = 73.48196394 / 1.651 * 1.651

       = 73.48196394 / 2.725801

       = 26.9579341779 Kg/M2

The BMI of the individual is approximately 27. This is considered overweight.

BMI weight ranges

  • Less than 18.5 = Underweight
  • Between 18.5 - 24.9 = Healthy Weight
  • Between 25 - 29.9 = Overweight
  • Over 30 = Obese